[Den, Mirko, Suresh]
We were investigating why there is no correlation between MC1 osem signals and seismic motion. During this we noticed a recurrence of this old problem of MC1_LR sensor being dead. I went and pressed down the chip holders where the AA filters used to sit and which now hold the jumper wire. The board is large and flexible it is quite likely some solder joint is broken on the MC1_LR path on this board.
The signal came back to life and is okay now. But it can break off again any time.
Since the MC1 LRSEN channel is not wasn't working, my input matrix diagonalization hasn't worked today wasn't working. So I decided to fix it somehow.
I went to the rack and traced the signal: first at the LEMO monitor on the whitening card, secondly at the 4-pin LEMO cable which goes into the AA chassis.
The signal existed at the input to the AA chassis but not in the screen. So I pressed the jumper wire (used to be AA filter) down for the channel corresponding to the MC1 LRSEN channel.
It now has come back and looks like the other sensors. As you can see from this plot and Joe's entry from a couple weeks ago, this channel has been dead since May 17th.
993190663 = free swinging ringdown restarted again